OpenStudy (anonymous):

f(x)=-4x-2 find f(2)

myininaya (myininaya):

So replace x with 2 then use order of operations to simplify and evaluate

OpenStudy (anonymous):

so x-1

OpenStudy (anonymous):

x=1

myininaya (myininaya):

no

myininaya (myininaya):

wherever you see an x you replace it with 2

myininaya (myininaya):

f(x)=-4x-2 f(2)=-4(2)-2 <---like this

OpenStudy (anonymous):

ok

myininaya (myininaya):

Now use order of operations to simplify/evaluate f(2)

myininaya (myininaya):

-4(2)-2

OpenStudy (anonymous):

-8

OpenStudy (anonymous):

-10

myininaya (myininaya):

-4(2)=-8 yes! And -8-2=-10 yes!

OpenStudy (anonymous):

thank